Your answer is partially correct. Try again h potential difference Vi 0.881 KV enters the gap between two parallel plates having separation d 23.6 mm and potential difference V plate is at the lower potential. Neglect fringing and assume that the electrons velocity vector is perpendicular to the electric feld vector allows the electron to travel in a straight line in the gap? o - 140 V. The lower between the plates. In unit-vector notation, what uniform magnetic neid ^i 1.337264???k ) Units Number t Click if you would like to Show Work for this questioni oen Shon rk
Answer #1

Energy, E = 0.5 mv^2 = q V

Velocity, v = sqrt( 2qV/m) = sqrt(2 x 1.6 x 10^-19 x 0.881 x 1000/ 9.1 x 10^-31) = 1.76 x 10^7 m/s

E = V'/ d = 140/ 0.0236 = 5932.20 N/C

B = E/v = 5932.20/(1.76 x 10^7) = -0.337 mT

Comment in case any doubt, please rate my answer....